OpenStudy (anonymous):

Aluminium chloride can be prepared by passing dry chlorine or hydrogen chloride gas over heated aluminium metal. The equation for the chlorine reaction is: 2Al(s) + 3Cl2(g) ==> 2AlCl3(s) How much aluminium is needed to make 89 g of aluminium chloride? (Relative atomic masses: Al = 27.0, Cl = 35.5, H =1) [rmc-29]

OpenStudy (anonymous):

\[89g AlCl_3 (\frac{1 mole AlCl_3}{223.5 g AlCl_3})(\frac{2 mole Al}{1 mole AlCl_3})(\frac{27.0 g Al}{1 mole Al})=21.5 g Al\] So you will need 21.5 grams of Al inorder to produce the 89g of AlCl3 that you needed.